OpenStudy (anonymous):

Find the angle between the vectors u and v if u = 3i +2j and v = 4i - 2j. Round answer to two decimal places.

OpenStudy (turingtest):

The dot product is\[\overrightarrow{u}*\overrightarrow{v}=|u||v|\cos\theta\]where |u| and |v| are the magnitudes of u and v, respectively. So we can solve this for theta...

OpenStudy (turingtest):

Can you do a dot product? Can you find the magnitude of the vectors u and v?

OpenStudy (anonymous):

yes i can, i just don't know how i would do it with the i and j. do i just leave them out?

OpenStudy (turingtest):

\[i*i=1\]\[j*j=1\]\[i*j=0\]so that is what will happen in the dot product. That way you wind up multiplying the individual components only.

OpenStudy (anonymous):

ohh i see! see i use the formula cos\[\cos \Theta= (u)(v)/ \left| u \right|\left| v \right|\]

OpenStudy (turingtest):

right, where \[(u)(v)\]is the dot product so inverse cosine both sides solves for theta

OpenStudy (anonymous):

i know how to do the dot product thing, just to get the denominator part how?

OpenStudy (turingtest):

The magnitude of a vector is essentially the length of the vector. So considering that each vector is made up of components at right angles to each other, we can use the Pythagorean theorem and treat the length of the vector as the hypotenuse.

OpenStudy (turingtest):

therefor\[|u|=\sqrt{(u_i)^2+(u_j)^2}\]where u_i and u_j are the components. We can do similarly for v.

OpenStudy (turingtest):

so what is\[|u|\]in your problem?

OpenStudy (anonymous):

u=3i+2j

OpenStudy (turingtest):

That is the vector itself. Use the formula above for the magnitude. Treat each component as legs of a right triangle and find the hypotenuse.

OpenStudy (turingtest):

look at the components:\[u_i=3\]\[u_j=2\]now use the formula above to find the magnitude.

OpenStudy (anonymous):

\[\sqrt{13}\]

OpenStudy (anonymous):

and v would be\[\sqrt{20}\]

OpenStudy (turingtest):

good :) so what is\[|v|\]???

OpenStudy (turingtest):

good

OpenStudy (turingtest):

now you can use\[\overrightarrow{u}*\overrightarrow{v}=|u||v|\cos\theta\]and solve for the angle.

OpenStudy (turingtest):

what do you have for\[\overrightarrow{u}*\overrightarrow{v}\]???

OpenStudy (anonymous):

i got 12-4=8

OpenStudy (turingtest):

right So we are ready to go:\[\overrightarrow{u}*\overrightarrow{v}=|u||v|\cos\theta\]solve for theta...

OpenStudy (anonymous):

so 8/ (\[\sqrt{13}\]\[\sqrt{20}\]

OpenStudy (anonymous):

inverse cosine

OpenStudy (anonymous):

theta equals 60.3

OpenStudy (turingtest):

yup! in degrees sounds about right... Nice job

OpenStudy (anonymous):

Thanks soo much!! :) that is one of the options! Thanks for all your help!

OpenStudy (turingtest):

anytime :)